OpenStudy (anonymous):

Two angles in a triangle have radian measure π/9 and π/2. What is the radian measure of the third angle?

OpenStudy (anonymous):

\[\pi-\frac{\pi}{9}-\frac{\pi}{2}\] as, \[2\pi^c=360^o\]